### v1

# Dynamic Post-Hoc Neural Ensemblers

Abstract
Ensemble methods are known for enhancing the accuracy and robustness of machine learning models by combining multiple base learners. However, standard approaches like greedy or random ensembles often fall short, as they assume a constant weight across samples for the ensemble members. This can limit expressiveness and hinder performance when aggregating the ensemble predictions. In this study, we explore employing neural networks as ensemble methods, emphasizing the significance of dynamic ensembling to leverage diverse model predictions adaptively. Motivated by the risk of learning low-diversity ensembles, we propose regularizing the model by randomly dropping base model predictions during the training. We demonstrate this approach lower bounds the diversity within the ensemble, reducing overfitting and improving generalization capabilities. Our experiments showcase that the dynamic neural ensemblers yield competitive results compared to strong baselines in computer vision, natural language processing, and tabular data.
1 Introduction
Ensembling machine learning models is a well-established practice among practitioners and researchers, primarily due to its enhanced predictive performance over single-model predictions. Ensembles are favored for their superior accuracy and ability to provide calibrated uncertainty estimates and increased robustness against covariate shifts (Lakshminarayanan et al., 2017). Combined with their relative simplicity, these properties make ensembling the method of choice for many applications, such as medical imaging and autonomous driving, where reliability is paramount.
Despite these advantages, the process of selecting post-hoc models that are both accurate and diverse remains a challenging combinatorial problem, especially as the pool of candidate models grows. Commonly used heuristics, particularly in the context of tabular data, such as greedy selection (Caruana et al., 2004) and various weighting schemes, attempt to optimize ensemble performance based on metrics evaluated on a held-out validation set or through cross-validation. However, these methods face significant limitations. Specifically, the selection of models to include in the ensemble and the determination of optimal ensembling strategies (e.g., stacking weights) are critical decisions that, if not carefully managed, can lead to overfitting on the validation data. Although neural networks are good candidates for generating ensembling weights, few studies rely on them as a post-hoc ensembling approach. We believe this happens primarily due to a lack of ensemble-related inductive biases that provide regularization.
In this work, we introduce a novel approach to post-hoc ensembling using neural networks. Our proposed Neural Ensembler dynamically generates the weights for each base model in the ensemble on a per-instance basis, a.k.a dynamical ensemble selection (Ko et al., 2008). To mitigate the risk of overfitting the validation set, we introduce a regularization technique inspired by the inductive biases inherent to the ensembling task. Specifically, we propose randomly dropping base models during training, inspired by previous work on DropOut in Deep Learning (Srivastava et al., 2014).
In summary, our contributions are as follows:
  1. 1.
    We propose a simple yet effective post-hoc ensembling method based on a neural network that dynamically ensembles base models.
  2. 2.
    To prevent the formation of low-diversity ensembles, we introduce a regularization technique that involves randomly dropping base model predictions during training. We demonstrate theoretically that this lower bounds the diversity of the generated ensemble. Additionally, we ablate its effect through various experiments.
  3. 3.
    Through extensive experiments, we show that Neural Ensemblers consistently select competitive ensembles across a wide range of data modalities, including tabular data (for both classification and regression), computer vision, and natural language processing.
To promote reproducibility, we have made our code publicly available in the following anonymous repository  [^1] https://github.com/machinelearningnuremberg/NeuralEnsemblers/. We hope that our codebase, along with the diverse set of benchmarks used in our experiments, will serve as a valuable resource for the development and evaluation of future post-hoc ensembling methods.
2 Background and Motivation
Post-Hoc ensembling uses set of fitted base models $\{z_{1},...,z_{M}\}$ such that every model outputs $z_{m}(x):\mathbb{R}^{D}\rightarrow\mathbb{R}$. These outputs are combined by a stacking ensembler $f(z(x);\theta):=f(z_{1}(x),...,z_{M}(x);\theta):\mathbb{R}^{M}\rightarrow\mathbb{R}$, where $z(x)=[z_{1}(x),...,z_{M}(x)]$ is the concatination of the base models predictions. While the base models are estimated using a training set $\mathcal{D}_{\mathrm{Train}}$, the ensembler’s parameters $\theta$ are typically obtained by minimizing a loss function on a validation set $\mathcal{D}_{\mathrm{Val}}$ such that:
$\theta\in\operatorname*{arg\,min}_{\theta}\;\sum_{(x,y)\in\mathcal{D}_{\mathrm{Val}}}\mathcal{L}(f(z(x);\theta),y).$ (1)
In the general case, this objective function can be optimized using gradient-free optimization method such as evolutionary algorithms (Purucker & Beel, 2023b) or greedy search (Caruana et al., 2004) together with a linear combination $\theta\in\mathbb{R}^{M}$ of the model outputs:
$f(z(x);\theta)=\sum_{m}\theta_{m}z_{m}(x).$ (2)
Additionally, if we constraint the ensembler weights such that $\forall_{i}\theta_{i}\in\mathbb{R}_{+}$ and $\sum_{i}\theta_{i}=1$ and assume probabilistic base models $z_{m}(x)=p(y|x,m)$, then we can interpret Equation 2 as:
$p(y|x)=\sum_{i}p(y|x,m)p(m),$ (3)
which is referred to as Bayesian Model Average, and uses $\theta_{m}=p(m)$. In the general case, the probabilistic ensembler $p(y|x)=p(y|z_{1}(x),...,z_{M}(x),\beta)$ is a stacker model parametrized by $\beta$.
2.1 Motivating Dynamic Ensembling
Figure 1: Wrong Models Per Samples Across Meta-Datasets. Every dark cell represents data instances where a model’s prediction is wrong. Different models fail on different instances, therefore, only instance-specific dynamic ensembles are optimal.
We motivate in this Section the need for dynamic ensembling by analyzing base models’ predictions in real data taken from our experimental metadatasets. In reality, Equation 3 does not specify the distribution $p(m)$. Generally, it is safe to assume that the performance associated with an ensembler $f(z_{m}(x),\theta)$ is optimal if we dynamically select the optimal aggregation $\theta_{m}(x)=p(m|m)$ on a per-data point basis, instead of a static $\theta$. To motivate this observation, we selected four datasets from different modalities: TabRepo (Tabular data (Salinas & Erickson, 2023)), QuickTune (Computer Vision (Arango et al., 2024)), FTC (NLP-Section D.1) and NasBench 201 (NAS for Computer Vision (Dong & Yang, 2020)). Then, we compute the per-sample error for 100 models in 100 samples. We report the results in Figure 1, indicating failed predictions with dark colors. We observe that models make different errors across samples, demonstrating the lack of optimality for static ensembling weights.
3 Neural Ensemblers
Figure 2: Architecture of Neural Ensemblers (classification). The stacking mode uses a single MLP shared across base model class predictions. It outputs the logit per class, used for computing the final probability via SoftMax. In Model Averaging mode, it generates the unnormalized weights for every model, which are normalized with SoftMax.
We use neural networks as ensemblers by training on the base model predictions $z(x)=[z_{1}(x);...;z_{M}(x)]$ as input. For regression, $z(x):\mathbb{R}^{D}\rightarrow\mathbb{R}^{M}$ outputs thebase models’ point predictions given by $x\in\mathbb{R}^{D}$, while for classification $z(x;c):\mathbb{R}^{D}\rightarrow[0,1]^{M}$ returns the probabilities predicted by the base models for class $c$. In our discussion we consider two functional modes for the ensemblers: as network outputting weights for model averaging or as a direct stacking model that outputs the prediction. In stacking mode for regression, we aggregate the base model point predictions using a neural network to estimate the final prediction $\hat{y}=\phi(z;\beta)$, where $\beta$ are the network parameters. In the model-averaging mode, the Neural Ensembler outputs the weights $\theta_{m}(z;\beta)$ to combine the model predictions as in Equation 4.
$\hat{y}=\sum_{m}\theta_{m}({z};\beta)\cdot z_{m}(x).$ (4)
Regardless of the functional mode, the Neural Ensembler has a different output $\hat{y}$ for regression and classification. In regression, the output $\hat{y}$ is a point estimation of the mean for a normal distribution such that $p(y|x;\beta)=\mathcal{N}(\hat{y},\sigma)$. For classification, the input is the probabilistic prediction of the base models per class $z_{m}(x;c)=p(y=c|x,m)$, while the output is a categorical distribution $\hat{y}=p(y=c|z(x),\beta)$. We optimize $\beta$ by minimizing the negative log-likelihood over the validation dataset $\mathcal{D}_{\mathrm{Val}}$ as:
$\min_{\beta}\mathcal{L}(\beta;D_{\mathrm{Val}})=\min_{\beta}\sum_{(x,y)\in\mathcal{D}_{\mathrm{Val}}}-\mathrm{log}\ p(y|x;\beta).$ (5)
3.1 Architecture
We discuss the architectural implementation of the Neural Ensembler for the classification case, which we show in Figure 2. For the stacking mode, we use an MLP that outputs the logit $\phi(z(x;c);\beta):\mathbb{R}^{M}\rightarrow[0,1]$ of each class $c$. Specifically, the network receives as inputs the base models’ predictions $z(x,c)$ for the class $c$ and outputs the corresponding predicted logit for this class, i.e. $\phi(z;c)$. The model predictions per class are fed independently, enabling sharing the network parameters $\beta$. Subsequently, we compute the probability $p(y=c|x)=\frac{\exp^{\phi(z;c)}}{\sum_{c^{\prime}}\exp^{\phi(z;c^{\prime})}}$, with $\mathrm{MLP}(z(x;c);\beta)$. In regression, the final prediction is the output $\phi(z(x);\beta)$.
For model averaging mode, we use a novel architecture based on a Deep Set (Zaheer et al., 2017) embedding of the base models predictions. We compute the dynamic weights $\theta_{m}(z;\beta)=\frac{\exp f_{m}(z;\beta)}{\sum_{m^{\prime}}f_{m^{\prime}}(z;\beta)}$, where the unnormalized weight per model $f_{m}(z;\beta)$ is determined via two MLPs. The first one $\text{MLP}_{1}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{H}$ embeds the predictions per class $z(x,c^{\prime})$ into a latent dimension of size $H$, whereas the second network $\text{MLP}_{2}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{H}$ aggregates the embeddings and outputs the unnormalized weights, as shown in Equation 6. Notice that the Neural Ensemblers’ input dimension and number of parameters are independent of the number of classes, due to our proposed parameter-sharing mechanism.
$f(z;m,\beta)=\mathrm{MLP}_{2}\left(\sum_{c^{\prime}}\mathrm{MLP}_{1}\left(z(x;c^{\prime});\beta_{1}\right);\beta_{2}\right).$ (6)
3.2 The Risk of Diversity Collapse
In this section, we elaborate on one risk that might arise during learning ensembles, as a base model might be highly correlated with the target in the validation set, when building ensembles. We dub this model a Preferred Base Model.
Definition 1 (Preferred Base Model).
Consider a target variable $y\in R$ and a set of uncorrelated base models predictions $\mathcal{Z}=\{z_{m}|z_{m}\in\mathrm{R},m=1,...,M\}$. $z_{p}$ is the Preferred Base Model if it has the highest sample correlation to the target, i.e. $\rho_{z_{p},y}\in[0,1],\rho_{z_{p},y}>\rho_{z_{m},y},\forall z_{m}\in\mathcal{Z}/\{z_{p}\}$.
Proposition 1.
If the correlation of the preferred model $\rho_{z_{p},y}\rightarrow 1$, then $\rho_{z_{i},y}\rightarrow 0,\forall z_{i}\in\mathcal{Z}/\{z_{p}\}$, given the conditions of Definition 1.
Sketch of Proof.
Given that the base models predictions are uncorrelated, their correlations $\rho_{z_{m},z_{p}}=0,\forall m\neq p$, then it holds that $-\sqrt{1-\rho_{z_{p},y}^{2}}\leq\rho_{z_{m},y}\leq\sqrt{1-\rho_{z_{p},y}^{2}},\forall z_{m}\in\mathcal{F}/\{z_{p}\}$. We provide more details in the Appendix A. ∎
On the other hand, an important aspect when building ensembles is guaranteeing diversity among the models (Wood et al., 2023; Jeffares et al., 2024). This has motivated some approaches to explicitly account for diversity when searching the ensemble configuration  (Shen et al., 2022; Purucker et al., 2023). A common way to measure diversity is the ambiguity (Krogh & Vedelsby, 1994), which can be derived after decomposing the loss function (Jeffares et al., 2024). Unfortunately, even simple ensembles risk obtaining low diversity although their base models are uncorrelated. As we will shortly show, this happens especially when there is a preferred base model.
Proposition 2 (Diversity Collapse).
If the correlation of the preferred model is $\rho_{z_{p},y}\rightarrow 1$ in an ensemble with prediction $\bar{z}=\sum_{m}\theta_{m}z_{m}$, then the ensemble diversity $\alpha\rightarrow 0$, where $\alpha:=\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$, i.e. $\lim_{\rho_{z_{p},y}\to 1}\alpha=0$.
We give a detailed proof in the Appendix A. However, we can build the intuition after noticing that given Proposition 1, $\lim_{\rho_{z_{p},y}\to 1}\bar{z}=z_{p}$.
3.3 Base Models’ DropOut
Input: Base model predictions $\{z_{1}(x),...,z_{M}(x)\}$, validation data $\mathcal{D}_{\mathrm{Val}}$, probability of retaining $\gamma$, mode $\in\{\mathrm{Stacking,Averaging}\}$.
Output: Neural Ensembler’s parameters $\beta$
1
2Initialize randomly parameters $\beta$ ;
3
4while $done$ do
5
6      Sample masking vector $r\in\mathbb{R}^{M},r_{i}\sim\mathrm{Ber}(\gamma)$;
7
8      Mask base models predictions $z_{\mathrm{drop}}(x)=r\odot z(x)$ ;
9
[…]
11             Compute predictions $\hat{y}=\phi\left(\frac{1}{\gamma}z_{\mathrm{drop}};\beta\right)$ ;
12      else
13
14            Compute weights $\theta\left(\frac{1}{\gamma}z_{\mathrm{drop}};m,\beta,r\right)$ using Equation 7 ;
15
16            Compute predictions $\hat{y}$ using Equation 4 ;
17
18       end if
19      Update neural ensembles parameters $\beta$ using $\nabla\mathcal{L}(\beta;\mathcal{D}_{\mathrm{Val}})$
20 end while
21
22return $\beta$;
23
Algorithm 1 Training Algorithm for Neural Ensemblers with Base Models’ DropOut
Using Neural Ensemblers tackles the need for dynamical ensembling. Moreover, it gives additional expressivity associated with neural networks. However, there is also a risk of overfitting and diversity collapse. As we showed in Section 3.2, this happens when there is a preferred model on which the ensembler mainly relies while neglecting other base model predictions. Although it might effectively decrease the validation loss (Equation 5), it does not necessarily generalize to test samples. Inspired by previous work (Srivastava et al., 2014), we propose to drop some base models during training forward passes. Intuitively, this forces the ensembler to rely on different base models to perform the predictions, instead of merely using the preferred base model(s).
Formally, we mask the inputs such as $r_{m}\cdot z_{m}(x)$, where $r_{m}\sim\mathrm{Ber}(\gamma)$, where $\mathrm{Ber}(\gamma)$ is the Bernoulli distribution with parameter $\gamma$ with represents the probability of keeping the base model, while $\delta=1-\gamma$ represents the DropOut rate. We also mask the weights when using model averaging:
$\theta(z;m,\beta,r)=\frac{r_{m}\cdot\exp{f_{m}}({z};\beta)}{\sum_{m^{\prime}}r_{m^{\prime}}\cdot\exp{f_{m^{\prime}}}({z};\beta)}.$ (7)
As DropOut changes the scale of the inputs, the weight scaling rule should be applied during inference by multiplying the dropped variables by the retention probability $\gamma$. Alternatively, we can scale the variables during training by $\frac{1}{\gamma}$. In Algorithm 1, we detail how to train the Neural Ensemblers by dropping base model predictions. It has two modes, acting as a direct stacker or as a model averaging ensembler. We demonstrate that base models’ DropOut avoids diversity collapse by upper bounding the diversity as stated in Proposition 6, even for the simplest ensembling case.
Proposition 3 (Avoiding Diversity Collapse).
As the correlation of the preferred model $\rho_{p_{m},y}\rightarrow 1$, the diversity $\alpha\rightarrow 1-\gamma$, when using Base Models’ DropOut with probability of retaining $\gamma$.
Sketch of Proof. We want to compute $\lim_{\rho_{z_{p},y}\to 1}\alpha=\lim_{\rho_{z_{p},y}\to 1}\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$. By using $\bar{z}=\sum_{m}r_{m}\ \theta_{m}z_{m}$, and assuming, without loss of generality, that the predictions are standardized, we obtain $\mathbb{V}(r\cdot z_{m})=\gamma$. This lead as to $\lim_{\rho_{z_{p},y}\to 1}\alpha=1-\gamma$, after following a procedure similar to Proposition 2. We provide the complete proof in Appendix A.
4 Experiments and Results
4.1 Proof-of-Concept: Ensemble of Quadratic Functions
To further highlight the importance of the dynamical ensemble with diverse base models, we propose a simple regression problem with a third-degree polynomial with the ground truth function $y_{\mathrm{true}}=1.33x^{3}-0.77x^{2}-0.31x-1$, where $x\in\left[-1,1\right]$. We fit three second-degree polynomials on different subsets of the training dataset, to obtain three base functions $z_{1}(x),z_{2}(x),z_{3}(x)$. As ensemblers we consider i) a static model average $\sum_{i}\theta_{m}z_{m}(x)$, ii) a dynamic model stacker $\phi(z(x);\beta)$ and iii) model average $\sum_{m}\theta_{m}(z(x);\beta)\cdot z_{m}(x)$. In this experiment, $\phi(\cdot)$ and $\theta_{m}(\cdot)$ are two-layer MLPs with 10 hidden neurons with parameters $\beta$ and trained on the validation data. Figure 3 shows the specific data, base functions, and learned ensemblers. The rightmost plot shows that the dynamic model averaging (MSE=$0.0101$) and stacker (MSE=$0.0055$), by expressing more complex functions, better model the ground truth than the static ensemble (MSE=$0.1662$).
Figure 3: Intuition Example for Dynamical Ensembling. When considering three base models $z_{1}(x),z_{2}(x),z_{3}(x)$, the dynamical ensemblers $\phi(\cdot)$ and $\theta_{m}(\cdot)$ depending on $x$ achieve better performance than a static model average, as observable in the right-most figure.
4.2 Experimental Setup
Meta-Datasets.
In our experiments, we utilize four meta-datasets with pre-computed predictions, which allows us to simulate ensembles without the need to fit models. These meta-datasets cover diverse data modalities, including Computer Vision, Tabular Data, and Natural Language Processing. Additionally, we evaluate the method on datasets without pre-computed predictions to assess the performance of ensembling methods that require model fitting. Table 1 reports the main information related to these datasets. Particularly for Nasbench, we created 2 versions by subsampling 100 and 1000 models. The metadataset for Finetuning Text Classifiers (FTC) was generated by ourselves to evaluate ensembling techniques on text classification tasks by finetuning language models such as GPT2 (Radford et al., 2019), Bert (Devlin et al., 2018) and Bart (Lewis et al., 2019). We also generate a set of fitted Scikit-Learn Pipelines on classification datasets. In this case, we stored the pipeline in memory, allowing us to evaluate our method in practical scenarios where the user has fitted models instead of predictions. We detailed information about the creation of these two metadatasets in Appendix D. Information about each dataset lies in the respective referred work under the column Task Information in Table 1.
Table 1: Metadatasets Information.
Meta-Dataset Modality Task Information No. Datasets Avg. Samples for Validation Avg. Samples for Test Avg. Models per Dataset
Nasbench (100) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 100
Nasbench (1K) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 1K
QuickTune (Micro) Vision Finetuning, Classification (Arango et al., 2024) 30 160 160 255
QuickTune (Mini) Vision Finetuning, Classification (Arango et al., 2024) 30 1088 1088 203
FTC Language Finetuning, Classification, Section D.1 6 39751 29957 105
TabRepo Clas. Tabular Classification  (Salinas & Erickson, 2023) 83 1134 126 1530
TabRepo Reg. Tabular Regression  (Salinas & Erickson, 2023) 17 3054 3397 1530
Sk-Learn Pipelines. Tabular Classification, Section D.2 69 1514 1514 500
Baselines.
We compare the Neural Ensemblers (NE) with other common and competitive ensemble approaches. 1) Single best selects the best model according to the validation metric; 2) Random chooses randomly $N=50$ models to ensemble, 3) Top-N ensembles the best $N$ models according to the validation metric; 4) Greedy creates an ensemble with $N=50$ models by iterative selecting the one that improves the metric as proposed by previous work (Caruana et al., 2004); 5) Quick builds the ensemble with 50 models by adding model subsequently only if they strictly improve the metric; 6) CMAES (Purucker & Beel, 2023b) uses an evolutive strategy with a post-processing method for ensembling, 7) Model Average (MA) computed the sum of the predictions with constant weights as in Equation 3. We also compare to methods that perform ensemble search iteratively via Bayesian Optimization such as 8) DivBO (Shen et al., 2022), and 9) Ensemble Optimization (EO) (Levesque et al., 2016). Finally, we report results by using common ML models as stackers, such as 10) SVM, 11) Random Forest, 12) Gradient Boosting, and 13) Logistic/Linear Regression. We used the default configurations provided by Sckit-learn (Pedregosa et al., 2011) for these stackers. The input to the models is the concatenation of all the base models’ predictions. We concatenated the probabilistic predictions from all the classes in the classification tasks. This sometimes produced a large dimensional input space. Finally, we include models from Dynamic Ensemble Search (DES) literature that are in the DESlib library (Cruz et al., 2020) such as KNOP (Cavalin et al., 2013), KNORAE (Ko et al., 2008) and MetaDES (Cruz et al., 2015).
Neural Ensemblers’ Setup.
We train the neural networks for $10000$ update steps, with a batch size of $2048$. If the GPU memory is not enough for fitting the network because of the number of base models, or the number of classes, we decreased the batch size to $256$. Additionally, we used the Adam optimizer and a network with four layers, $32$ neurons, and a probability of keeping base models $\gamma=0.25$, or alternatively a DropOut rate $\delta=0.75$. Notice that the architecture of the ensemblers slightly varies depending on the mode (Stacking or Model Average). For the Stacking mode, we use an MLP with four layers and $32$ neurons with ReLU activations. For MA mode, we use two MLPs as in Equation 6: 1) $\textrm{MLP}_{1}$ has $3$ layers with $32$ neurons, while 2) $\textrm{MLP}_{2}$ has one layer with the same number of neurons. Although changing some of these hyperparameters might improve the performance, we keep this setup constant for all the experiments, after checking that the Neural Ensemblers perform well in a subset of the Quick-Tune metadataset (extended version).
4.3 Research Questions and Associated Experiments
Table 2: Average Normalized Error.
FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Class (AUC)
Single-Best 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Random 1.3377±0.2771 0.7283±0.0752 0.7491±0.2480 6.6791±3.4638 4.7284±2.9463 1.4917±1.6980 1.7301±1.8127
Top5 0.9511±0.0364 0.6979±0.0375 0.6296±0.1382 0.6828±0.3450 0.8030±0.2909 0.9998±0.1233 0.9271±0.2160
Top50 1.1012±0.1722 0.6347±0.0395 0.5650±0.1587 1.0662±0.9342 1.0721±0.4671 0.9800±0.1773 0.9297±0.2272
Quick 0.9494±0.0371 0.6524±0.0436 0.5787±0.1510 0.7575±0.2924 0.7879±0.2623 0.9869±0.1667 0.9054±0.2232
Greedy 0.9494±0.0374 0.7400±0.1131 1.0000±0.0000 0.9863±0.4286 0.9297±0.1435 0.9891±0.1693 0.9090±0.2197
CMAES 0.9489±0.0392 0.6401±0.0343 0.5797±0.1575 1.0319±0.5000 0.9086±0.1121 0.9935±0.1953 1.1878±1.1457
Random Forest 0.9513±0.0359 0.6649±0.0427 0.6891±0.3039 1.4738±1.3510 1.2530±0.4875 1.0041±0.2330 1.0924±0.6284
Gradient Boosting 1.0097±0.1033 1.2941±0.5094 1.2037±0.3528 0.8514±0.5003 1.6121±1.7023 1.0452±0.3808 1.0663±0.4884
SVM 0.9453±0.0383 0.6571±0.0483 0.7015±0.3067 1.1921±0.8266 1.4579±0.6233 0.9585±0.2160 1.4701±1.3486
Linear 0.9609±0.0347 0.7891±0.1978 0.7782±0.1941 0.7333±0.4457 0.9291±0.3580 0.9776±0.2844 1.0329±0.4022
MA 1.0709±0.0845 0.6381±0.0349 0.5610±0.1490 1.1548±0.8465 1.2173±0.6107 1.0917±1.0135 0.9977±0.2278
DivBO 1.0155±0.1452 0.6915±0.0536 0.9120±0.1524 1.3935±1.4316 1.0635±0.7587 1.0908±1.0104 1.0899±1.0297
EO 1.0208±0.1159 0.6365±0.0445 0.5704±0.1619 1.0185±0.6464 1.0367±0.4394 1.0851±1.0136 0.9377±0.2310
NE-Stack (Ours) 0.9491±0.0451 0.6331±0.0378 0.5836±0.1592 0.6104±0.3656 0.7545±0.2960 1.0440±0.3309 1.0035±0.5295
NE-MA (Ours) 0.9527±0.0402 0.6307±0.0363 0.5621±0.1483 0.8297±0.4974 0.8236±0.2240 0.9592±0.2144 0.9028±0.2157
RQ 1: Can Neural Ensemblers outperform other common and competitive ensembling methods across data modalities?
Figure 4: Results on Scikit-Learn Pipelines.
Experimental Protocol. To answer this question, we compare the neural ensembles in stacking and averaging mode to the baselines across all the metadatasets. We run every ensembling method three times for every dataset. In all the methods we use the validation data for fitting the ensemble, while we report the results on the test split. Specifically, we report the average across datasets of two metrics: negative log-likelihood (NLL) and error. For the tabular classification, we compute the ROC-AUC. As these metrics vary for every dataset, we normalize metrics by dividing them by the single-best metric. Therefore, a method with a normalized metric below one is improving on top of using the single best base model. We report the standard deviation across the experiments per dataset and highlight in bold the best method.
Results. The results reported in Table 2 and Table 3 show that our proposed neural networks are competitive post-hoc ensemblers. In general, we observe that the Neural Ensemblers variants obtain either the best (in bold) or second best (italic) performance across almost all meta-datasets and metrics. Noteworthily, the greedy approach is very competitive, especially for the FTC and TR-Class. This is coherent with previous work supporting greedy ensembling as a robust method for tabular data (Erickson et al., 2020). We hypothesize that dynamic ensembling contributes partially to the strong results for the Neural Ensemblers. However, the expressivity gained is not enough, because it can lead to overfitting. To understand this, we compare to Dynamic Ensemble Selection (DES) methods. Specifically, we use KNOP, MetaDES, and KNORAE, and evaluate all methods in Scikit-learn Pipelines metadataset, as we can easily access the fitted models. We report the results of the test split in Figure 4, where we distinguish among four types of models to facilitate the reading: Neural, DES, Stacking and Constant. We can see that Neural Ensemblers are the most competitive approaches, especially on stacking mode. Additionally, we report the metrics on the validation split in Figure 8 (Appendix E), where we observe that some dynamic ensemble approaches such as Gradient Boosting (GBT), Random Forest (RF) and KNORAE exhibit overfitting, while Neural Enemblers are more robust in this sense.
Table 3: Average Normalized NLL.
FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single-Best 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Random 1.5450±0.5289 0.6591±0.2480 0.7570±0.2900 6.8911±3.1781 5.8577±3.2546 1.7225±1.9645 1.8319±2.1395
Top5 0.8406±0.0723 0.6659±0.1726 0.6789±0.3049 1.5449±1.8358 1.1496±0.3684 1.0307±0.5732 0.9939±0.0517
Top50 0.8250±0.1139 0.5849±0.2039 0.6487±0.3152 3.3068±2.6197 3.0618±2.2960 1.0929±1.0198 1.0327±0.2032
Quick 0.7273±0.0765 0.5957±0.1940 0.6497±0.3030 1.1976±1.1032 0.9747±0.2082 0.9860±0.2201 1.0211±0.1405
Greedy 0.6943±0.0732 0.5785±0.1972 1.0000±0.0000 0.9025±0.2378 0.9093±0.1017 0.9665±0.0926 1.0149±0.1140
CMAES 1.2356±0.5295 1.0000±0.0000 1.0000±0.0000 4.1728±2.8724 4.6474±3.0180 1.3487±1.3390 1.0281±0.1977
Random Forest 0.7496±0.0940 0.8961±0.3159 0.9340±0.4262 3.7033±2.8145 2.2938±2.2068 1.2655±0.4692 1.0030±0.0871
Gradient Boosting 0.7159±0.1529 1.7288±1.2623 1.2764±0.4787 1.9373±1.2839 2.6193±2.3159 1.4288±1.2083 1.0498±0.2128
SVM 0.7990±0.0909 0.7744±0.2967 0.9358±0.5706 5.4377±3.3807 4.0019±3.6601 1.3884±1.4276 2.7975±3.0219
Linear 0.7555±0.0898 0.7400±0.2827 0.8071±0.2206 1.3960±1.2334 1.1031±0.7038 1.1976±1.1024 3.1488±3.2813
MA 0.9067±0.1809 0.5970±0.2034 0.6530±0.3028 4.7921±3.0780 4.0168±2.8560 1.4724±1.9401 1.3342±1.3515
DivBO 0.7695±0.1195 0.7307±0.3061 0.7125±0.3982 1.2251±1.0293 0.9430±0.2036 1.0023±0.3411 1.0247±0.1473
EO 0.7535±0.1156 0.5801±0.2051 0.6911±0.2875 1.3702±1.6389 0.9649±0.2980 1.0979±1.0289 1.0183±0.0993
NE-Stack (Ours) 0.7562±0.1836 0.5278±0.2127 0.6336±0.3456 0.7486±0.6831 0.6769±0.2612 1.3268±0.7498 1.2379±0.4083
NE-MA (Ours) 0.6952±0.0730 0.5822±0.2147 0.6522±0.3131 1.0177±0.5151 0.9166±0.0936 1.0515±1.0003 0.9579±0.0777
RQ 2: Do Neural Ensemblers need a strong group of base models, i.e. found using Bayesian Optimization?
Experimental Protocol. Practitioners use some methods such as greedy ensembling as post-hoc ensemblers, i.e., they consider a set of models selected by a search algorithm such as Bayesian Optimization as base learners. DivBO enhances the Bayesian Optimization by accounting for the diversity in the ensemble in the acquisition function. We run experiments to understand whether the Neural Ensemblers’ performance depends on a strong subset of $50$ base models selected by DivBO, and whether it can help other methods. We conduct additional experiments by randomly selecting $50$ models to understand the impact and significance of merely using a smaller set of base models. We normalize the base of the metric on the single-best base model from the complete set contained in the respective dataset.
Results. We report in Table 4 the results with the two selection methods (random and DivBO) using a subset of common baselines, where we normalize using the metric of the single-best from the whole set of models. We limit the number of baselines for brevity, and the extended version includes more baselines in Table 15. We also include results in separate tables for the two selection methods (Appendix E). We can compare directly with the results in Table 3. We observe that reducing the number of base models with DivBO negatively affects the performance of the Neural Ensemblers. Surprisingly, randomly selecting the subset of base models improves the results in two metadatasets (TR-Class and NB-1000). We hypothesize that decreasing the number of base models is beneficial for these metadatasets. With over $1000$ base models available, the likelihood of identifying a preferred model and overfitting the validation data increases in these metadatasets. Naturally, decreasing the number of base models can also be detrimental for the Neural Ensemblers, as this happens for some metadatasets such as TR-Reg and QT-Micro. In contrast to the Neural Ensemblers, selecting a subset of strong models with DivBO improves the performance for some baselines such as Model Averaging (MA) or TopK ($K=25$). In other words, it works as a preprocessing method for these ensembling approaches. Overall, the results in Tables 4 and 15 demonstrate that Neural Ensemblers do not need a strong group of base models to achieve competitive results.
Table 4: Average Normalized NLL with a Subset of Base Models.
Selector FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single - 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Single DivBO 1.0000±0.0000 1.0000±0.0000 0.8707±0.3094 1.7584±2.0556 1.1846±0.2507 1.1033±0.9951 1.0039±0.0424
Random DivBO 0.9305±0.3286 0.6538±0.2123 0.9724±0.0478 1.1962±1.0189 0.9717±0.1919 1.0107±0.3431 1.0302±0.1250
Top25 DivBO 0.7617±0.1136 0.5564±0.1961 0.9762±0.0413 1.1631±0.9823 0.9431±0.2035 1.0023±0.3411 1.0247±0.1473
Quick DivBO 0.7235±0.0782 0.6137±0.1945 0.9646±0.0614 1.2427±1.1130 0.9544±0.2050 1.0014±0.3423 1.0400±0.1949
Greedy DivBO 0.7024±0.0720 0.6839±0.3003 0.9762±0.0413 1.1659±0.9789 0.9435±0.2029 1.0024±0.3410 1.0271±0.1531
MA DivBO 0.7245±0.0788 0.5712±0.2185 0.9678±0.0558 1.0559±0.7452 0.9501±0.1617 1.0068±0.4141 1.0237±0.1502
Single Random 1.0067±0.0164 1.0000±0.0000 0.9240±0.3504 1.2915±0.9952 1.1261±0.3134 1.0225±0.3353 1.1378±0.4641
Top25 Random 0.8397±0.1000 0.5848±0.1980 0.6526±0.3019 3.6553±2.7053 3.0436±2.1378 1.2599±1.5015 1.0611±0.2799
Quick Random 0.7305±0.0764 0.5958±0.1917 0.6656±0.2968 1.7769±2.1443 1.1646±0.3728 1.0797±1.0007 1.0151±0.1546
Greedy Random 0.7024±0.0720 0.5783±0.1857 0.6617±0.2839 1.6723±2.1446 0.9961±0.1290 1.0725±0.9978 1.0023±0.0961
MA Random 0.9069±0.1812 0.8677±0.2292 0.6698±0.2898 4.8593±3.1360 3.4575±2.6490 1.4759±1.9396 1.4286±1.7242
NE-Stack (Ours) DivBO 0.7715±0.2141 0.6204±0.2234 1.0000±0.0000 1.5040±1.9442 0.8329±0.2659 0.9729±0.3952 6.9453±3.4749
NE-MA (Ours) DivBO 0.7036±0.0698 0.5704±0.2345 1.0000±0.0000 1.1237±0.9964 0.9200±0.1966 1.0016±0.3407 1.0070±0.0977
NE-Stack (Ours) Random 0.7709±0.2204 0.7551±0.2493 0.6187±0.2950 0.8292±0.5466 0.8160±0.3852 0.9540±0.5077 4.2183±3.4808
NE-MA (Ours) Random 0.6972±0.0712 0.7911±0.2147 0.6650±0.2750 1.6877±2.1535 1.0903±0.2578 1.0674±0.9998 1.0277±0.1994
RQ 3: What is the impact of the DropOut regularization scheme?
Experimental Protocol. Firstly, we do a small experiment to observe how the DropOut rate affects the weights when using Neural Ensemblers in MA mode. For this, we use QT-Micro datasets, and save each model’s weights $\theta_{m}$ and error after training the neural network. Subsequently, to understand how much the base learners DropOut helps the Neural Ensemblers, we run an ablation by trying the following values for the DropOut rate $\delta\in\{0.0,0.1,\dots,0.9\}$. We compute the average NLL for three seeds per dataset and divide this value by the one obtained for $\delta=0.0$ in the same dataset. Therefore, we realize that a specific DropOut rate is improving over the default network without regularization if the normalized NLL is below 1.
Results. We show how the mean of the weights per model is related to the mean error of the models for different DropOut rates in Figure 5. When there is no DropOut some weights are close to one, i.e. they are preferred models. As we increase the value, many models with high weights decrease. If the rate is very high (e.g. $0.9$), we will have many models contributing to the ensemble, with weights different from zero. Our ablation study demonstrates that non-existing or high DropOut are detrimental to the Neural Ensembler performance in general. As shown in Figure 6, this behavior is consistent in all datasets and both modes, but TR-Reg metadataset on Stacking mode. In general, we observe that Neural Ensemblers obtain better performance when using base models’ DropOut.
Figure 5: Mean weights assigned to the base models decrease with DropOut rate. Every data point is a model. The errors and weights are the mean values across many datasets for every model.
Figure 6: Ablation of the DropOut rate.
5 Related Work
Ensembles for Tabular Data. For tabular data, ensembles are known to perform better than individual models (Sagi & Rokach, 2018; Salinas & Erickson, 2023).Therefore, ensembles are often used in real-world applications (Dong et al., 2020), to win competitions (Koren, 2009; Kaggle, 2024), and by automated machine learning (AutoML) systems as a modeling strategy (Purucker & Beel, 2023b; Purucker et al., 2023). Methods like Bagging (Breiman, 1996) or Boosting (Freund et al., 1996) are often used to boost the performance of individual models. In contrast, post-hoc ensembling (Shen et al., 2022; Purucker & Beel, 2023a) aggregates the predictions of an arbitrary set of fitted base models. Post-hoc ensembles are build by stacking (Wolpert, 1992; Van der Laan et al., 2007), ensemble selection (a.k.a. pruning) (Caruana et al., 2004; Tsoumakas et al., 2009), dynamic ensemble selection (Ko et al., 2008; Britto Jr et al., 2014), or through a systematic search for an optimal ensemble (Levesque et al., 2016; Shen et al., 2022; Poduval et al., 2024).
Ensembles for Deep Learning. Ensembles of neural networks (Hansen & Salamon, 1990; Krogh & Vedelsby, 1994; Dietterich, 2000) have gained significant attention in deep learning research, both for their performance-boosting capabilities and their effectiveness in uncertainty estimation. Various strategies for building ensembles exist, with deep ensembles (Lakshminarayanan et al., 2017) being the most popular one, which involve independently training multiple initializations of the same network. Extensive empirical studies (Ovadia et al., 2019; Gustafsson et al., 2020) have shown that deep ensembles outperform other approaches for uncertainty estimation, such as Bayesian neural networks (Blundell et al., 2015; Gal & Ghahramani, 2016; Welling & Teh, 2011).
Mixture-of-Experts. Our idea of generating ensemble base model weights is closely connected to the mixture-of-experts (MoE) (Jacobs et al., 1991; Jordan & Jacobs, 1993; Shazeer et al., 2017), where one network is trained with specialized sub-modules that are activated based on the input data. Batch ensembles (Wen et al., 2020; Wenzel et al., 2020) are also closely related, as they aim to capture diverse model behaviors by simultaneously training multiple models with shared weights but different random projections. In contrast to MoE and batch ensembles, our Neural Ensemblers can ensemble any (black-box) model and are not restricted to gradient-based approaches. Ours additionally considers a regularization strategy that leads to provably more diverse ensembles.
Dynamic Ensemble Selection. Our Neural Ensembler is highly related to dynamic ensemble selection. Both dynamically aggregate the predictions of base models per instance (Cavalin et al., 2013; Ko et al., 2008). Traditional dynamic ensemble selection methods aggregate the most competent base models by paring heuristics to measure competence with clustering, nearest-neighbor-based, or traditional tabular algorithms (like naive Bayes) as meta-models (Cruz et al., 2018; 2020). In contrast, we use an end-to-end trained neural network to select and weight the base models per instance.
6 Conclusions
In this work, we tackled the challenge of post-hoc ensemble selection and the associated risk of overfitting on the validation set. We introduced the Neural Ensembler, a neural network that dynamically assigns weights to base models on a per-instance basis. To reduce overfitting, we proposed a regularization technique that randomly drops base models during training, which we theoretically showed enhances ensemble diversity. Our empirical results demonstrated that Neural Ensemblers consistently form competitive ensembles across diverse data modalities, including tabular data (classification and regression), computer vision, and natural language processing. In the future, we aim to explore in-context learning (Brown et al., 2020), where a pretrained Neural Ensembler could generate base model weights at test time, using their predictions as contextual input. We discuss broader impact and limitations in Appendix B.
Reproducibility Statement
All code and datasets used in this paper are publicly available at https://github.com/machinelearningnuremberg/NeuralEnsemblers/. We provide the code for running all the experiments, including baselines on all meta-datasets. We also provide instructions for the environment setup to ensure reproducibility. We will deanonymize the link upon acceptance. We further provide all proofs in Appendix A. Moreover, we ensured reproducibility by testing our approach on many meta-datasets and across several domains, for which details can be found in Appendix D. We also compared our method to many baselines commonly used in ensembling, described in Section 4.3. Lastly, we detail additional results and extended data tables in Appendix E.
Acknowledgments
The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project v101be. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. This research was partially supported by the following sources: TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828 and 499552394 - SFB 1597; the European Research Council (ERC) Consolidator Grant “Deep Learning 2.0” (grant no. 101045765). Frank Hutter acknowledges financial support by the Hector Foundation. The authors acknowledge support from ELLIS and ELIZA. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the ERC. Neither the European Union nor the ERC can be held responsible for them.
Appendix
Appendix A Proofs
Definition 2 (Preferred Base Model).
Consider a target variable $y\in R$ and a set of uncorrelated base models predictions $\mathcal{Z}=\{z_{m}|z_{m}\in\mathrm{R},m=1,...,M\}$. $z_{p}$ is the Preferred Base Model if it has the highest sample correlation with respect the target, i.e. $\rho_{z_{p},y}\in[0,1],\rho_{z_{p},y}>\rho_{z_{m},y},\forall z_{m}\in\mathcal{Z}/\{z_{p}\}$.
Proposition 4.
If the correlation of the preferred model $\rho_{z_{p},y}\rightarrow 1$, then $\rho_{z_{i},y}\rightarrow 0,\forall z_{i}\in\mathcal{Z}/\{z_{k}\}$, given the conditions of Definition 1.
Proof.
Let $\rho_{z_{p},z_{m}},\rho_{z_{m},y},\rho_{z_{p},y}$ be the correlations among the preferred model $z_{p}$, a non-preferred model $z_{m}\neq z_{m}$ and the target $y$, then the correlation matrix is given by,
$C=\begin{pmatrix}1&\rho_{z_{p},z_{m}}&\rho_{z_{p},y}\\ \rho_{z_{p},z_{m}}&1&\rho_{z_{m},y}\\ \rho_{z_{p},y}&\rho_{z_{m},y}&1\\ \end{pmatrix}$ (8)
The determinant of the correlation matrix is given by,
$\mathrm{Det}(C)$ $=1-\rho_{z_{p},z_{m}}^{2}-\rho_{z_{p},y}^{2}-\rho_{z_{m},y}^{2}+2\rho_{z_{p},z_{m}}\rho_{z_{p},y}\rho_{z_{m},y}$ (9)
$\mathrm{Det}(C)$ $=1-\rho_{z_{p},y}^{2}-\rho_{z_{m},y}^{2}$ (10)
$\mathrm{Det}(C)$ $=1-\rho_{z_{p},y}^{2}-\rho_{z_{m},y}^{2}\geq 0$ (11)
$\rho_{z_{m},y}$ $\leq\sqrt{1-\rho_{z_{p},y}^{2}},$ (12)
where we used the fact that $\rho_{z_{p},z_{m}}=0$, in Equation 10 and that the correlation matrix should be semidefinite in Equation 11, thus its determinant should hold $\mathrm{Det}(C)\geq 0$. This means that $\epsilon=\sqrt{1-\rho_{z_{p},y}^{2}}$ is an upper bound for the absolute value of the correlation $\rho_{z_{m},y}$. The limit of the upper bound $\epsilon$ as $\rho_{z_{p},y}\to 1$ is 0 and thus $\rho_{z_{m},y}=0$,
$\lim_{\rho_{z_{p},y}\to 1}$ $\epsilon=\lim_{\rho_{z_{p},y}\to 1}\sqrt{1-\rho_{z_{p},y}^{2}}=0$ (13)
$\lim_{\rho_{z_{p},y}\to 1}$ $\rho_{z_{m},y}=0.$ (14)
Since $\rho_{z_{m},y}$ is bounded by $-\epsilon\leq\rho_{z_{m},y}\leq\epsilon$.
Proposition 5 (Diversity Collapse).
If the correlation of the preferred model is $\rho_{z_{p},y}\rightarrow 1$ in an ensemble with prediction $\bar{z}=\sum_{m}\theta_{m}z_{m}$, then the ensemble diversity $\alpha\rightarrow 0$, where $\alpha:=\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$.
Proof.
Firstly, we show that under conditions of Definition 1, $\theta_{m}=\hat{\rho}_{z_{m},y}$, where $\hat{\rho}_{z_{m},y}$ is the sample correlation. Assume that we have a matrix of evaluations $X\in\mathrm{R}^{N\times M}$, with components $x_{n,m}$ indicating the prediction of the $m$-th model for the $n$-th instance. If $Y\in\mathbb{R}^{N\times 1}$ is the ground-truth, then the ensemble weights $\Theta\in\mathbb{R}^{M\times 1}$ can be computed with closed form solution of the objective $\min_{\Theta}||Y-X\Theta||^{2}$ as: $\Theta=(X^{T}X)^{-1}X^{T}Y$.
Without loss of generality, we assume that the random variables are standardized. Given that every row of $X$ have samples from uncorrelated random variables, then $X^{T}X\approx I$, denoting $I$ as the identity matrix, thus
$\Theta=X^{T}Y$ (15)
Every component of $\Theta$ can expressed as summatorias $\theta_{m}=\sum_{n}x_{n,m}y_{n}$. Note that $x_{n,m}$ is a standardized sample (i.e. mean 0, variance 1) from the random variable $z_{m}$, thus it equals the sample correlation $\theta_{m}=\sum_{n}x_{n,m}y_{n}=\hat{\rho}_{z_{m},y}$.
Now we develop the formal demonstration of Proposition 2. We note to the right side a hint to the criteria that was applied to derive every equation.
$\lim_{\rho_{z_{p},y}\to 1}\alpha=\lim_{\rho_{z_{p},y}\to 1}$ $\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$ Definition of $\alpha$. (16)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\mathbb{E}\left[(z_{m}-\bar{z})^{2}\right]$ Properties of expectations. (17)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(\mathbb{E}[z_{m}-\bar{z}]^{2}+\mathbb{V}[z_{m}-\bar{z}]\right)$ Expectation of a squared variable. (18)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\mathbb{V}[z_{m}-\bar{z}]$ Means of $z_{m}$ and $\bar{z}$ are 0. (19)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(\mathbb{V}[z_{m}]-\mathbb{V}[\bar{z}]\right)$ Property of sum of variances. (20)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\mathbb{V}[\bar{z}]\right)$ Variance of $z_{m}$ is 1. (21)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\mathbb{V}\left[\sum_{m^{\prime}}\theta_{m^{\prime}}z_{m^{\prime}}\right]\right)$ Definition of $\bar{z}$. (22)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\mathbb{V}\left[z_{m^{\prime}}\right]\right)$ Properties of variance. (23)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\right)$ Variance of $z_{m}$ is 1. (24)
$=\sum_{m}\lim_{\rho_{z_{p},y}\to 1}$ $\theta_{m}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\right)$ Property of sum of limits. (25)
$=\sum_{m}\lim_{\rho_{z_{p},y}\to 1}$ $\theta_{m}\cdot\lim_{\rho_{z_{p},y}\to 1}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\right)$ Product-sum property of limits. (26)
$=\left[\sum_{m}\lim_{\rho_{z_{p},y}\to 1}\rho_{z_{m},y}\right]$ $\cdot\left[\lim_{\rho_{z_{p},y}\to 1}\left(1-\sum_{m^{\prime}}\rho_{z_{m^{\prime}},y}^{2}\right)\right]$ Given that $\theta_{m}=\hat{\rho}_{z_{m},y}\approx\rho_{z_{m},y}$. (27)
=1=1$\cdot$= 1 ⋅ $\left[\lim_{\rho_{z_{p},y}\to 1}\left(1-\sum_{m^{\prime}}\rho_{z_{m^{\prime}},y}^{2}\right)\right]$ Given Proposition 1. (28)
$=1-$ $\sum_{m^{\prime}}\lim_{\rho_{z_{p},y}\to 1}\rho_{z_{m^{\prime}},y}^{2}$ Properties of limits. (29)
$\lim_{\rho_{z_{p},y}\to 1}$ $\alpha=0$ Given Proposition 1. (30)
Lemma 1.
When using dropout, it still holds that $\theta_{m}=\hat{\rho}_{z_{m},y}.$
Proof.
$\min_{\Theta}\mathbb{E}_{R\sim\mathrm{Ber(\gamma)}}||Y-(R*X)\Theta||^{2}$ (31)
$=\min_{\Theta}||Y-\gamma X\Theta||^{2}+\gamma(1-\gamma)||\Gamma\Theta||^{2},$ (32)
where $\Gamma=(\mathrm{diag(X^{T}X)})^{1/2}$. If we use the uncorrelation of the base models, we obtain:
$=\min_{\Theta}||Y-\gamma X\Theta||^{2}+\gamma(1-\gamma)||\Theta||^{2}$ (33)
$=\min_{\Theta}||Y-\tilde{X}\Theta||^{2}+\gamma(1-\gamma)||\Theta||^{2},$ (34)
which is equivalent to ridge regression, with closed form solution:
$\Theta=$ $\left({\tilde{X}}^{T}{\tilde{X}}+\gamma(1-\gamma)I\right)^{-1}{\tilde{X}}^{T}Y$ (35)
$=$ $\left([\gamma^{2}+\gamma(1-\gamma)]I\right)^{-1}{\tilde{X}}^{T}Y$ (36)
$=$ $\left(\gamma I\right)^{-1}{\tilde{X}}^{T}Y$ (37)
$=$ $\frac{1}{\gamma}{\tilde{X}}^{T}Y$ (38)
$=$ $X^{T}Y.$ (39)
Proposition 6 (Avoiding Diversity Collapse).
As the correlation of the preferred model $\rho_{p_{m},y}\rightarrow 1$, the diversity $\alpha\rightarrow 1-\gamma$, when using Base Models’ DropOut with probability of retaining $\gamma$.
Proof.
We follow a similar procedure as for Proposition 2, by considering $\bar{z}=\sum_{m}r_{m}\ \theta_{m}z_{m}$. We demonstrate that $\mathbb{V}(r\cdot z_{m})=\gamma$, given that $r\sim\mathrm{Bernoulli}(\gamma)$.
$\mathbb{V}(r_{m}\cdot z_{m})$ $=\mathbb{V}(r_{m})\cdot\mathbb{V}(z_{m})+\mathbb{V}(z_{m})\cdot\mathbb{E}(r_{m})^{2}+\mathbb{V}(r_{m})\cdot\mathbb{E}(z_{m})^{2}$ (40)
$\mathbb{V}(r_{m}\cdot z_{m})$ $=\mathbb{V}(r_{m})+\mathbb{E}(r_{m})^{2}$ (41)
$\mathbb{V}(r_{m}\cdot z_{m})$ $=\gamma(1-\gamma)+\gamma^{2}$ (42)
$\mathbb{V}(r_{m}\cdot z_{m})$ $=\gamma.$ (43)
Then, we evaluate the variance of the ensemble using DropOut $\mathbb{V}(\bar{z})$:
$\mathbb{V}(\bar{z})=\mathbb{V}{\left(\sum_{m}r_{m}\cdot\theta_{m}\cdot z_{m}\right)}$ (44)
$\mathbb{V}(\bar{z})=\sum_{m}\mathbb{V}{\left(r_{m}\cdot\theta_{m}\cdot z_{m}\right)}$ (45)
$\mathbb{V}(\bar{z})=\sum_{m}\theta^{2}_{m}\mathbb{V}{\left(r_{m}\cdot z_{m}\right)}$ (46)
$\mathbb{V}(\bar{z})=\gamma\sum_{m}\theta^{2}_{m}.$ (47)
Applying Equation 47 into Equation 22, we obtain:
$\lim_{\rho_{z_{p},y}\to 1}\alpha$ $=\lim_{\rho_{z_{p},y}\to 1}\mathbb{E}\left[\sum_{m}\gamma_{m}\cdot\theta_{m}(z_{m}-\bar{z})^{2}\right]$ (48)
$=\lim_{\rho_{z_{p},y}\to 1}\sum_{m}\theta_{m}\left(1-\gamma\sum_{m}\theta^{2}_{m}\right)$ (49)
$=\lim_{\rho_{z_{p},y}\to 1}\left(1-\gamma\sum_{m^{\prime}}\rho_{z_{m^{\prime}},y}^{2}\right)$ (50)
$=1-\gamma\cdot\lim_{\rho_{z_{p},y}\to 1}\left(\sum_{m^{\prime}}\rho_{z_{m^{\prime}},y}^{2}\right)$ (51)
$\lim_{\rho_{z_{p},y}\to 1}\alpha$ $=1-\gamma.$ (52)
Appendix B Limitations and Broader Impact.
While our proposed method offers several advantages for post-hoc ensemble selection, it is important to recognize its limitations. Unlike simpler ensembling heuristics, our approach requires tuning multiple training and architectural hyperparameters. Although we employed a fixed set of hyperparameters across all modalities and tasks in our experiments, this robustness may not generalize to all new tasks. In such cases, hyperparameter optimization may be necessary to achieve optimal performance. However, this could also enhance the results presented in this paper. Our approach is highly versatile and can be seamlessly integrated into a wide variety of ensemble-based learning systems, significantly enhancing their predictive capabilities. Because our method is agnostic to both the task and modality, we do not expect any inherent negative societal impacts. Instead, its effects will largely depend on how it is applied within different contexts and domains, making its societal implications contingent on the specific use case.
Appendix C Related Work Addendum
Ensemble Search via Bayesian Optimization. Ensembles of models with different hyperparameters can be built using Bayesian optimization by iteratively swapping a model inside an ensemble with another one that maximizes the expected improvement (Levesque et al., 2016). DivBO (Shen et al., 2022) and subsequent work (Poduval et al., 2024) combine the ensemble’s performance and diversity as a measure for expected improvement. Besides Bayesian Optimization, an evolutionary search can find robust ensembles of deep learning models (Zaidi et al., 2021). Although these approaches find optimal ensembles, they can overfit the validation data used for fitting if run for many iterations.
Appendix D Details On Metadatasets
We provide general information about the datasets in Table 5.
Table 5: Metadatasets Information
Meta-Dataset Modality Task Information No. Datasets Avg. Samples for Validation Avg. Samples for Test Avg. Models per Dataset
Nasbench (100) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 100
Nasbench (1K) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 1K
QuickTune (Micro) Vision Finetuning, Classification (Arango et al., 2024) 30 160 160 255
QuickTune (Mini) Vision Finetuning, Classification (Arango et al., 2024) 30 1088 1088 203
FTC Language Finetuning, Classification, Section D.1 6 39751 29957 105
TabRepo Clas. Tabular Classification  (Salinas & Erickson, 2023) 83 1134 126 1530
TabRepo Reg. Tabular Regression  (Salinas & Erickson, 2023) 17 3054 3397 1530
Sk-Learn Pipelines. Tabular Classification, Section D.2 69 1514 1514 500
D.1 Finetuning Text Classifiers (FTC) Metadataset
Table 6: Search Space.
Hyperparameter Values
Model GPT2, Bert-Large, Albert-Large, Bart-Large, T5-Large
Learning Rate 0.00001, 0.0001, 0.0005, 0.001, 0.005
LoRA Rank 8, 16, 32, 64, 128
Search Space
It comprises three hyperparameters: model, learning rate and LoRA rank (Hu et al., 2022). We consider consider five models: 1) GPT2, 124M parameters; (Radford et al., 2019); 2) Bert-Large, 336M ; (Devlin et al., 2018); 3) Bart-Large 400 M, parameters  (Lewis et al., 2019); 4) Albert-Large, 17M parameters  (Lan et al., 2019); and 5) T5-Large, 770 M parameters  (Raffel et al., 2020). For the other two hyperparameters we also consider five different values as specified in Table 6.
Datasets
The metadataset contains predictions of models finetuned in five metadatasets for text classification: 1) IMDB (Maas et al., 2011); 2) Tweet (Maggie, 2020), 3) News (Zhang et al., 2015), 4) DBpedia (Zhang et al., 2015), 5) SST2 (Socher et al., 2013) and 6) SetFit (Tunstall et al., 2021). For every dataset, we create two versions. The first one is trained with the complete training data, while the second version is only with a subset of 10 % of the samples. All the datasets are for text classification beween 2 to 14 classes, including diverse domains such as movies, reviews, news, tweets, and text entailment data. We provide further information about the datasets in Table 7.
Figure 7: Mean error across datasets.
Metadataset Creation and Composition
We created the dataset by finetuning every model to the train split and, subsequently, saving their predictions on the validation and test split. The validation split corresponds to 20 % of the available train data. As for some datasets (SST-2 and SetFit) the test data is not completely provided by the creators, or it has hidden labels, we obtain it by using 20 % of the remaining test data. Specifically The models are fitted up to 5 epochs, using 1 GPU A100 with batch size equal to 2 and no LoRA dropout. We vary only the model, learning rate, and LoRA rank, while keeping the default hyperparameters in the Trainer object from the Transformers Library (version 4.41.0). Although we evaluate the models in a grid, some runs yielded out-of-memory errors for some configurations. In total, the metadataset contains 1134 evaluated configurations, representing around 3800 GPU hours. Additionally, we report information about the metadataset in Table 7.
Table 7: Metadataset Information
Dataset # Classes # Train Samples # Val. Samples # Test Samples # Confs (100%) # Confs. (10%)
IMDB (Maas et al., 2011) 2 20,000 5,000 25,000 125 125
Tweet (Maggie, 2020) 3 27,485 5,497 3,534 100 100
News (Zhang et al., 2015) 4 96,000 24,000 7,600 99 120
DBpedia (Zhang et al., 2015) 14 448,000 112,000 70,000 25 65
SST-2  (Socher et al., 2013) 2 43,103 13,470 10,776 125 125
SetFit  (Tunstall et al., 2021) 3 393,116 78,541 62,833 25 100
Hyperparameter Importance
We explore the importance of two hyperparameters, learning rate, and LoRA rank, by plotting the mean error as a heatmap in Figure 7. The error corresponds to the average across different models and datasets. The learning rate is an important hyperparameter, while increasing the LoRA rank does not affect the performance significantly in low learning rates. This behaviour is interesting, as it showcases that a small rank is enough for a successful finetuning in this context.
D.2 Scikit learn Pipelines
Search Space
Our primary motivation is to investigate the ensembling of automated machine learning pipelines to enhance performance across various classification tasks. To effectively study ensembling methods and benchmark different strategies, we require a diverse set of pipelines. Therefore, we construct a comprehensive search space inspired by the TPOT library (Olson et al., 2016), encompassing a wide range of preprocessors, feature selectors, and classifiers. The pipelines are structured in three stages—preprocessor, feature selector, and classifier—which allows us to systematically explore numerous configurations. This extensive and diverse search space enables us to examine the impact of ensembling on a variety of models and serves as a robust benchmark for evaluating different ensembling techniques. Detailed descriptions of the components and their hyperparameters are provided in Tables 8, 9, and 10.
Table 8: Classifiers and their hyperparameters used in the TPOT search space.
Classifier Hyperparameters
sklearn.naive_bayes.GaussianNB None
sklearn.naive_bayes.BernoulliNB
alpha (float, [1e-3, 100.0], default=50.0)
fit_prior (categorical, {True, False})
sklearn.naive_bayes.MultinomialNB
alpha (float, [1e-3, 100.0], default=50.0)
fit_prior (categorical, {True, False})
sklearn.tree.DecisionTreeClassifier
criterion (categorical, {’gini’, ’entropy’})
max_depth (int, [1, 10], default=5)
min_samples_split (int, [2, 20], default=11)
min_samples_leaf (int, [1, 20], default=11)
sklearn.ensemble.ExtraTreesClassifier
n_estimators (constant, 100)
criterion (categorical, {’gini’, ’entropy’})
max_features (float, [0.05, 1.0], default=0.525)
min_samples_split (int, [2, 20], default=11)
min_samples_leaf (int, [1, 20], default=11)
bootstrap (categorical, {True, False})
sklearn.ensemble.RandomForestClassifier
n_estimators (constant, 100)
criterion (categorical, {’gini’, ’entropy’})
max_features (float, [0.05, 1.0], default=0.525)
min_samples_split (int, [2, 20], default=11)
min_samples_leaf (int, [1, 20], default=11)
bootstrap (categorical, {True, False})
sklearn.ensemble.GradientBoostingClassifier
n_estimators (constant, 100)
learning_rate (float, [1e-3, 1.0], default=0.5)
max_depth (int, [1, 10], default=5)
min_samples_split (int, [2, 20], default=11)
min_samples_leaf (int, [1, 20], default=11)
subsample (float, [0.05, 1.0], default=0.525)
max_features (float, [0.05, 1.0], default=0.525)
sklearn.neighbors.KNeighborsClassifier
n_neighbors (int, [1, 100], default=50)
weights (categorical, {’uniform’, ’distance’})
p (categorical, {1, 2})
sklearn.linear_model.LogisticRegression
penalty (categorical, {’l1’, ’l2’})
C (float, [1e-4, 25.0], default=12.525)
dual (categorical, {True, False})
solver (constant, ’liblinear’)
xgboost.XGBClassifier
n_estimators (constant, 100)
max_depth (int, [1, 10], default=5)
learning_rate (float, [1e-3, 1.0], default=0.5)
subsample (float, [0.05, 1.0], default=0.525)
min_child_weight (int, [1, 20], default=11)
n_jobs (constant, 1)
verbosity (constant, 0)
sklearn.linear_model.SGDClassifier
loss (categorical, {’log_loss’, ’modified_huber’})
penalty (categorical, {’elasticnet’})
alpha (float, [0.0, 0.01], default=0.005)
learning_rate (categorical, {’invscaling’, ’constant’})
fit_intercept (categorical, {True, False})
l1_ratio (float, [0.0, 1.0], default=0.5)
eta0 (float, [0.01, 1.0], default=0.505)
power_t (float, [0.0, 100.0], default=50.0)
sklearn.neural_network.MLPClassifier
alpha (float, [1e-4, 0.1], default=0.05)
learning_rate_init (float, [0.0, 1.0], default=0.5)
Table 9: Preprocessors and their hyperparameters used in the TPOT search space.
Preprocessor Hyperparameters
None None
sklearn.preprocessing.Binarizer
threshold (float, [0.0, 1.0], default=0.5)
sklearn.decomposition.FastICA
tol (float, [0.0, 1.0], default=0.0)
sklearn.cluster.FeatureAgglomeration
linkage (categorical, {’ward’, ’complete’, ’average’})
metric (categorical, {’euclidean’, ’l1’, ’l2’, ’manhattan’, ’cosine’})
sklearn.preprocessing.MaxAbsScaler None
sklearn.preprocessing.MinMaxScaler None
sklearn.preprocessing.Normalizer
norm (categorical, {’l1’, ’l2’, ’max’})
sklearn.kernel_approximation.Nystroem
kernel (categorical, {’rbf’, ’cosine’, ’chi2’, ’laplacian’, ’polynomial’,
      ’poly’, ’linear’, ’additive_chi2’, ’sigmoid’})
gamma (float, [0.0, 1.0], default=0.5)
n_components (int, [1, 10], default=5)
sklearn.decomposition.PCA
svd_solver (categorical, {’randomized’})
iterated_power (int, [1, 10], default=5)
sklearn.preprocessing.PolynomialFeatures
degree (constant, 2)
include_bias (categorical, {False})
interaction_only (categorical, {False})
sklearn.kernel_approximation.RBFSampler
gamma (float, [0.0, 1.0], default=0.5)
sklearn.preprocessing.RobustScaler None
sklearn.preprocessing.StandardScaler None
tpot.builtins.ZeroCount None
tpot.builtins.OneHotEncoder
minimum_fraction (float, [0.05, 0.25], default=0.15)
sparse (categorical, {False})
threshold (constant, 10)
Table 10: Feature selectors and their hyperparameters used in the TPOT search space.
Selector Hyperparameters
None None
sklearn.feature_selection.SelectFwe
alpha (float, [0.0, 0.05], default=0.025)
sklearn.feature_selection.SelectPercentile
percentile (int, [1, 100], default=50)
sklearn.feature_selection.VarianceThreshold
threshold (float, [0.0001, 0.2], default=0.1)
sklearn.feature_selection.RFE
step (float, [0.05, 1.0], default=0.525)
estimator (categorical, {’sklearn.ensemble.ExtraTreesClassifier’})
Estimator Hyperparameters:
   n_estimators (constant, 100)
   criterion (categorical, {’gini’, ’entropy’})
   max_features (float, [0.05, 1.0], default=0.525)
sklearn.feature_selection.SelectFromModel
threshold (float, [0.0, 1.0], default=0.5)
estimator (categorical, {’sklearn.ensemble.ExtraTreesClassifier’})
Estimator Hyperparameters:
   n_estimators (constant, 100)
   criterion (categorical, {’gini’, ’entropy’})
   max_features (float, [0.05, 1.0], default=0.525)
Datasets We utilized the OpenML Curated Classification benchmark suite 2018 (OpenML-CC18) (Bischl et al., 2019) as the foundation for our meta-dataset. OpenML-CC18 comprises $72$ diverse classification datasets carefully selected to represent a wide spectrum of real-world problems, varying in size, dimensionality, number of classes, and domains. This selection ensures comprehensive coverage across various types of classification tasks, providing a robust platform for evaluating the performance and generalizability of different ensembling approaches.
Metadataset Creation To construct our meta-dataset, we randomly selected $500$ pipeline configurations for each dataset from our comprehensive search space. Each pipeline execution was constrained to a maximum runtime of $15$ minutes. During this process, we had to exclude three datasets (connect-4, Devnagari-Script, Internet-Advertisements) due to excessive computational demands that exceeded our runtime constraints. For data preprocessing, we standardized the datasets by removing missing values and encoding categorical features. We intentionally left other preprocessing tasks to be handled autonomously by the pipelines themselves, allowing them to adapt to the specific characteristics of each dataset. This approach ensures that the pipelines can perform necessary transformations such as scaling, normalization, or feature engineering based on their internal configurations, which aligns with our objective of evaluating automated machine learning pipelines in a realistic setting.
Appendix E Additional Results
In this Section we report additional results from our experiments:
  • Average Ranking of baselines for the Negative Log-likelihood (Table 11) and Classification Errors 12.
  • Average NLL in Scikit-learn pipelines metadataset 8.
  • Average NLL using a subset of base models selected via DivBO 13, random 14 or both 15.
Table 11: Average Ranked NLL
FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Reg
Single-Best 14.0000±0.8944 14.0000±1.0000 12.8333±1.0408 6.8833±2.3219 7.3667±2.5049 7.5361±3.8531 7.7647±3.9096
Random 16.0000±0.0000 7.6667±4.0415 9.3333±0.5774 15.2500±0.8068 15.2333±0.7739 12.0000±3.9962 13.6176±1.8669
Top5 11.3333±2.5033 9.6667±1.1547 7.0000±1.0000 6.9000±1.9360 8.3000±1.9678 6.2470±3.3951 7.2353±4.2907
Top50 10.8333±1.7224 4.6667±1.1547 4.3333±0.5774 11.4000±1.2959 11.5333±1.5025 6.8554±3.4006 6.8824±3.7397
Quick 6.1667±1.6021 6.3333±3.0551 3.6667±2.5166 4.7333±2.0331 6.0333±2.5661 6.2048±2.8531 5.9412±3.2107
Greedy 3.5000±0.8367 4.0000±2.6458 12.8333±1.0408 3.5167±1.7786 4.5667±2.4238 5.9036±3.0268 6.5882±3.2607
CMAES 12.1667±4.4460 14.0000±1.0000 12.8333±1.0408 12.9667±2.6811 13.9667±1.3322 9.9277±3.8880 6.2353±2.4630
Random Forest 7.0000±2.7568 13.0000±2.6458 14.0000±2.6458 11.9833±1.2964 9.3000±3.8699 11.4217±5.1493 8.2353±4.8157
Gradient Boosting 3.1667±4.8339 14.6667±1.5275 13.5000±2.1794 9.7000±5.2203 10.1167±5.1890 11.4217±4.9463 8.3529±4.7558
SVM 10.6667±1.3663 10.0000±7.0000 11.3333±7.2342 13.9167±1.2532 11.6833±5.0198 10.8434±4.2641 11.7941±5.5707
Linear 7.6667±2.0656 11.0000±1.0000 10.0000±2.0000 6.8000±2.8816 6.5833±3.3479 10.3855±5.2587 14.3529±3.4989
MA 11.1667±2.8577 6.6667±1.1547 4.3333±2.3094 13.2167±1.0059 12.8167±1.6000 9.0241±4.0567 8.2941±5.1813
DivBO 7.1667±3.5449 9.0000±4.5826 5.3333±5.7735 4.9167±2.5123 5.1000±2.7114 6.6145±2.9377 6.7647±2.9054
EO 5.8333±3.7103 5.3333±1.5275 7.3333±1.5275 5.8833±2.2194 5.5000±2.5052 6.2169±2.6689 7.0000±2.5739
NE-Stack (Ours) 6.0000±4.7329 1.0000±0.0000 3.3333±4.0415 3.2000±3.4381 2.3333±2.0899 10.0241±5.8351 12.2941±3.5314
NE-MA (Ours) 3.3333±2.5033 5.0000±3.0000 4.0000±1.7321 4.7333±2.0331 5.5667±1.9945 5.3735±3.2674 4.6471±2.3702
Table 12: Average Ranked Error
FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Class (AUC)
Single-Best 12.1667±2.5626 15.0000±0.8660 13.8333±0.7638 9.8167±3.7908 9.1167±3.5591 8.9157±4.1480 9.3086±4.1030
Random 15.8333±0.4082 13.0000±0.0000 12.0000±1.0000 15.9667±0.1826 15.9667±0.1826 11.3253±5.0723 12.2716±4.3747
Top5 6.0833±4.2710 11.6667±2.0817 8.3333±2.0817 4.4000±2.8780 4.6167±2.9204 8.7048±3.6325 7.2099±4.0519
Top50 13.5000±2.2583 3.3333±1.5275 2.0000±1.0000 8.3000±4.1618 9.7333±2.6546 7.9880±4.1548 7.2222±3.6929
Quick 6.3333±4.1433 7.6667±1.1547 6.3333±2.3094 5.5833±3.5186 3.5500±2.3575 8.3253±3.1771 6.8148±3.8033
Greedy 5.0833±3.8264 11.3333±3.0551 13.8333±0.7638 9.0000±3.5012 6.9833±3.1961 8.0843±3.3467 6.9321±3.2984
CMAES 4.5833±2.2004 5.0000±3.6056 6.3333±2.0817 8.5833±3.1103 6.3167±2.1794 8.8313±3.7776 9.6605±4.5537
Random Forest 6.6667±4.4121 8.3333±4.6188 9.5000±4.4441 11.7000±3.6237 12.5000±1.9343 8.8313±4.0817 8.7778±4.5600
Gradient Boosting 10.5000±3.2711 15.6667±0.2887 14.5000±1.5000 8.4667±4.2729 12.4000±3.1139 9.1024±4.9840 9.0123±5.0812
SVM 3.5000±3.3317 8.0000±1.7321 8.6667±5.1316 10.1167±4.8382 13.3667±2.6682 7.0602±3.9296 10.5185±5.4906
Linear 7.8333±2.6394 11.6667±3.5119 11.3333±2.0817 5.7500±3.2450 7.5167±2.9870 7.7892±5.2179 9.0185±5.1705
MA 12.8333±2.4014 4.6667±0.5774 2.3333±1.5275 10.0667±3.5349 11.3667±2.9271 8.4277±4.2949 8.7531±3.7200
DivBO 9.4167±3.3229 10.6667±1.1547 12.6667±1.5275 10.0833±2.9916 7.3833±3.2262 8.1506±3.6599 8.4938±3.4228
EO 10.0000±4.6476 4.6667±3.2146 4.0000±2.0000 9.3667±2.8555 8.5000±2.6425 8.1988±4.2677 7.6852±2.8410
NE-Stack (Ours) 4.6667±3.0768 3.3333±1.1547 7.3333±0.5774 3.4500±2.3095 2.7167±2.2232 8.9639±5.4352 7.6173±5.5437
NE-MA (Ours) 7.0000±1.7889 2.0000±1.0000 3.0000±2.0000 5.3500±3.7374 3.9667±2.2967 7.3012±3.9900 6.7037±3.7855
Table 13: Average NLL for Subset of Base Models with DivBO
Selector FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single - 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Single DivBO 1.0000±0.0000 1.0000±0.0000 0.8707±0.3094 1.7584±2.0556 1.1846±0.2507 1.1033±0.9951 1.0039±0.0424
Random DivBO 0.9305±0.3286 0.6538±0.2123 0.9724±0.0478 1.1962±1.0189 0.9717±0.1919 1.0107±0.3431 1.0302±0.1250
Top25 DivBO 0.7617±0.1136 0.5564±0.1961 0.9762±0.0413 1.1631±0.9823 0.9431±0.2035 1.0023±0.3411 1.0247±0.1473
Quick DivBO 0.7235±0.0782 0.6137±0.1945 0.9646±0.0614 1.2427±1.1130 0.9544±0.2050 1.0014±0.3423 1.0400±0.1949
Greedy DivBO 0.7024±0.0720 0.6839±0.3003 0.9762±0.0413 1.1659±0.9789 0.9435±0.2029 1.0024±0.3410 1.0271±0.1531
MA DivBO 0.7245±0.0788 0.5712±0.2185 0.9678±0.0558 1.0559±0.7452 0.9501±0.1617 1.0068±0.4141 1.0237±0.1502
NE-Stack (Ours) DivBO 0.7715±0.2141 0.6204±0.2234 1.0000±0.0000 1.5040±1.9442 0.8329±0.2659 0.9729±0.3952 6.9453±3.4749
NE-MA (Ours) DivBO 0.7036±0.0698 0.5704±0.2345 1.0000±0.0000 1.1237±0.9964 0.9200±0.1966 1.0016±0.3407 1.0070±0.0977
Table 14: Average NLL for Subset of Base Models randomly
Seletor FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single - 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Single Random 1.0067±0.0164 1.0000±0.0000 0.9240±0.3504 1.2915±0.9952 1.1261±0.3134 1.0225±0.3353 1.1378±0.4641
Top25 Random 0.8397±0.1000 0.5848±0.1980 0.6526±0.3019 3.6553±2.7053 3.0436±2.1378 1.2599±1.5015 1.0611±0.2799
Quick Random 0.7305±0.0764 0.5958±0.1917 0.6656±0.2968 1.7769±2.1443 1.1646±0.3728 1.0797±1.0007 1.0151±0.1546
Greedy Random 0.7024±0.0720 0.5783±0.1857 0.6617±0.2839 1.6723±2.1446 0.9961±0.1290 1.0725±0.9978 1.0023±0.0961
MA Random 0.9069±0.1812 0.8677±0.2292 0.6698±0.2898 4.8593±3.1360 3.4575±2.6490 1.4759±1.9396 1.4286±1.7242
NE-Stack Random 0.7709±0.2204 0.7551±0.2493 0.6187±0.2950 0.8292±0.5466 0.8160±0.3852 0.9540±0.5077 4.2183±3.4808
NE-MA Random 0.6972±0.0712 0.7911±0.2147 0.6650±0.2750 1.6877±2.1535 1.0903±0.2578 1.0674±0.9998 1.0277±0.1994
Table 15: Average NLL for Subset of Base Models (more baselines)
Selector FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single DivBO 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Random DivBO 0.9305±0.3286 0.6538±0.2123 0.9767±0.0404 0.8220±0.3507 0.8814±0.2062 0.9748±0.1415 1.0287±0.1365
Top25 DivBO 0.7617±0.1136 0.5564±0.1961 0.9805±0.0339 0.7916±0.3245 0.8569±0.2178 0.9672±0.1381 1.0232±0.1550
Quick DivBO 0.7235±0.0782 0.6137±0.1945 0.9687±0.0542 0.8048±0.3271 0.8672±0.2195 0.9652±0.1353 1.0378±0.1959
Greedy DivBO 0.7024±0.0720 0.6839±0.3003 0.9805±0.0339 0.7910±0.3209 0.8572±0.2174 0.9673±0.1384 1.0256±0.1601
CMAES DivBO 0.8915±0.1759 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 0.9671±0.1179 1.0250±0.1595
Random Forest DivBO 0.7932±0.1194 0.9338±0.3436 1.1678±0.2906 1.6807±1.1160 1.5507±0.8449 1.3195±0.6428 1.0534±0.1277
Gradient Boosting DivBO 0.7908±0.1848 1.4011±0.5359 1.0000±0.0000 2.4553±1.1191 2.7220±1.3007 1.2106±0.5937 1.1246±0.4052
Linear DivBO 0.7433±0.0870 0.6471±0.2272 1.0248±0.0430 0.9690±0.4209 1.1132±0.4525 1.0814±1.0056 1.0316±0.1403
SVM DivBO 0.8312±0.0943 0.7406±0.2612 1.0786±0.1362 5.2463±3.4452 4.8194±3.5169 1.5156±1.5597 2.7777±2.9458
MA DivBO 0.7245±0.0788 0.5712±0.2185 0.9719±0.0486 0.7603±0.2876 0.8802±0.1741 0.9621±0.1306 1.0222±0.1570
Single Random 1.0067±0.0164 1.0000±0.0000 1.0540±0.0936 0.8958±0.3297 0.9786±0.2444 0.9902±0.1529 1.1314±0.4475
Top25 Random 0.8397±0.1000 0.5848±0.1980 0.7283±0.1534 2.7131±1.5510 2.5961±1.9153 1.2873±1.5773 1.0588±0.2762
Quick Random 0.7305±0.0764 0.5958±0.1917 0.7467±0.1420 0.9847±0.1836 0.9737±0.1340 1.0043±0.2497 1.0123±0.1529
Greedy Random 0.7024±0.0720 0.5783±0.1857 0.7478±0.1352 0.9171±0.0932 0.8960±0.0795 0.9762±0.1082 1.0002±0.1038
CMAES Random 1.2164±0.3851 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.1544±0.9890 0.9985±0.0954
RF Random 0.7505±0.0915 0.8325±0.2255 0.9850±0.0175 2.9838±1.7970 2.6001±1.8694 1.2006±0.4367 1.0058±0.0898
GBT Random 0.7235±0.1605 1.8377±1.4510 1.2177±0.3771 1.9521±1.2076 2.8683±1.9763 1.2329±0.6934 1.0470±0.1913
Linear Random 0.7541±0.0897 0.7400±0.2827 0.9103±0.1384 1.1776±0.3187 1.5498±0.8446 1.0417±1.0146 1.0408±0.1296
SVM Random 0.8010±0.0903 0.7767±0.3006 0.9821±0.3676 5.1993±3.4269 5.1080±3.2902 1.4479±1.4990 2.8328±3.0214
MA Random 0.9069±0.1812 0.8677±0.2292 0.7543±0.1312 4.1694±2.5844 3.0315±2.4578 1.5059±1.9654 1.4185±1.6833
NE-Stack Random 0.7709±0.2204 0.7551±0.2493 0.6882±0.1602 0.6925±0.4510 0.6854±0.2801 0.9387±0.4628 4.2065±3.4678
NE-MA Random 0.6972±0.0712 0.7911±0.2147 0.7535±0.1164 0.9291±0.0999 0.9191±0.0624 0.9922±0.2525 1.0259±0.2002
NE-Stack DivBO 0.7715±0.2141 0.6204±0.2234 1.0000±0.0000 1.0442±0.8691 0.7482±0.2254 0.9671±0.3746 6.9574±3.4368
NE-MA DivBO 0.7036±0.0698 0.5704±0.2345 1.0000±0.0000 0.7632±0.3319 0.8349±0.2075 0.9677±0.1455 1.0057±0.1121
Figure 8: Average Normalized NLL in Scikit-learn pipelines (Validation).